OpenStudy (geekfromthefutur):

Is square root of four hundred rational or irrational? Explain your reasoning.

OpenStudy (geekfromthefutur):

I think its irrational

OpenStudy (mathstudent55):

Do you know how to find the prime factors of 400?

OpenStudy (anonymous):

Square root of 400 is 20, since 20 x 20 is 400. And 20 is certainly rational since it can be expressed as a fraction using integers 20 = 20/1

OpenStudy (mathstudent55):

Now that you have the answer, let's figure out how you do it. Do you know how to find the prime factors of 400?

OpenStudy (geekfromthefutur):

??????????

OpenStudy (mathstudent55):

Do you know what prime numbers are?

OpenStudy (geekfromthefutur):

Yea

OpenStudy (mathstudent55):

Every integer greater than 1 is either a prime number or a product of prime numbers.

OpenStudy (anonymous):

I personally wouldn't bother with prime factors on thisone. When you see 400, you should think what squared equals 400? 2x2=4 and there are 2 0's so, try 20x20..hey it works. Faster than prime factorisation.

OpenStudy (mathstudent55):

For example, 2, 3, 5, 7, 11 are prime numbers. 6 is not a prime number since 6 = 2 * 3, a product of prime numbers.

OpenStudy (mathstudent55):

To find the prime factorization of a number, divide the number by each prime number starting with 2 as many times as possible. When it is no longer divisible by 2, move on to 3. Keep going until you get 1.

OpenStudy (geekfromthefutur):

Ok im getting it so far

OpenStudy (mathstudent55):

Here it is with 400. 400/2 = 200 200/2 = 100 100/2 = 50 50/2 = 25 25 is not divisible by 2, so we move to 3. 25 is also not divisible by 3, so we move to 5. 25/5 = 5 5/5 = 1 Since we finally got 1, the factorization is done. Now we see what it is. Now look at the prime numbers we divided by: 2, 2, 2, 2, 5, 5 Those numbers are the prime factorization of 400. In other words, 400 = 2 * 2 * 2 * 2 * 5 * 5 = 2^4 * 5^2

OpenStudy (mathstudent55):